A ball is thrown vertically upward with a speed of 20m/s when another ball is from a height 50m is released. So find out the time when they meet

Grade:9

1 Answers

Rohan
6 years ago
Let the position where ball meet be xAnd let the distance covered by first ball be x then the distance by the second ball is 50-xS=ut+1/2gt^2Initial velocity of second ball is zero,u=0S=1/2gt^2 gtTaking g=9.8m/s^2 gt50-x=1/2gt^2 equation 1X=ut-1/2gt^2 equation 2From 1 and 250-x= 20×t -(50-x)20t=50T=2.5 seconds They will meet after 2.5seconds